This Question: 2 pts 19 of 23 (18 complete) Weights (in grams) of randomly selected chocolate candies are represented in the table below. Find the range and standard deviation. 0.952 0.913 0.923 0.843 0.937 0.881 0913 0.912 0.953 0.947 0.919 Find the range. range-9 Find the standard deviation s-g (Round to four decimal places as neoded.)
Answer #1

Ans:

Minimum=0.843

Maximum=0.953

Range=maximum-minimum

=0.953-0.843=0.11

xi (xi-0.917545)^2
1 0.952 0.001187
2 0.913 0.000021
3 0.923 0.000030
4 0.843 0.005557
5 0.937 0.000378
6 0.881 0.001336
7 0.913 0.000021
8 0.912 0.000031
9 0.953 0.001257
10 0.947 0.000868
11 0.919 0.000002
Total= 10.093 0.010687
mean= 0.917545

n=11

sample standard deviation,s=sqrt(0.010687/(11-1))=0.03269

*(population standard deviation=sqrt(0.010687/11)=0.03117)